OpenStudy (anonymous):

Solve: -1/2 x + 2 = −x + 7 10 5/2 2/5 20

OpenStudy (***[isuru]***):

hi, \[- \frac{ 1 }{ 2 }x+ 2 = -x + 7\] \[x - \frac{ 1 }{ 2 }x = 7 -2\] \[\frac{ 1 }{ 2}x = 5\] \[x = 10\] hope this will help ya!!!